# A Comprehensive Survey on Knowledge Distillation

### v1

@@ VII Performance Comparison
ILD [484] (2022) ICL $GPT2_{large}$ 6.00 CrossFit 61.20 / 66.20 92.0%
LLM-R [485] (2023) ICL $LLaMA_{13B}$ 2.00 CommonsenseQA 48.80 / 49.60 98.0%
AICD [486] (2024) ICL GPT-3.5-Turbo SVAMP 51.70 / 20.70 250.0%
UniversalNER [479] (2023) IF GPT-3.5-Turbo UNIVERSAL NER 41.70 / 34.90 119.0% Link
LaMini-LM [480] (2023) IF $Alpaca_{7B}$ 9.00 OpenBookQA 34.00 / 43.20 79.0% Link
Lion [481] (2023) IF $GPT_{175B}$ 25.00 BBH 32.00 / 48.90 65.0% Link
Fine-tune-CoT [468] (2022) CoT $GPT_{175B}$ 26.00 SVAMP 30.33 / 64.67 47.0% Link
Li et al. [466] (2022) CoT $GPT_{175B}$ 58.00 CommonsenseQA 82.47 / 73.00 113.0%
Magister et al. [467] (2022) CoT $GPT_{175B}$ 16.00 StrategyQA 63.77 / 65.40 97.0%
MCC-KD [469] (2023) CoT GPT-3.5-Turbo SVAMP 68.66 / 75.14 91.0% Link
CSoTD [471] (2023) CoT $GPT_{175B}$ 135.00 CommonsenseQA 67.00 / 82.10 82.0%
SCOTT [472] (2023) CoT $GPT\,neox_{20B}$ 7.00 CommonsenseQA 74.70 / 60.40 124.0% Link
Distilling step-by-step [470] (2023) CoT $Palm_{540B}$ 700.00 SVAMP 58.40 / 72.30 81.0% Link
PaD [474] (2023) CoT GPT-3.5-Turbo SVAMP 36.70 / 57.80 64.0%
TDG [475] (2024) CoT GPT-3.5-Turbo MATH 6.81 / 3.88 175.0% Link
RevThink [476] (2024) CoT Gemini-1.5-Pro-001 30.00 CommonsenseQA 75.76 / 76.72 99.0%
In this work, existing methods are compared for two well-known tasks: classification and semantic segmentation segmentation, as almost all the proposed distillation methods evaluate their results on these two tasks. For classification, the CIFAR-100 dataset, and for semantic segmentation, the Cityscapes dataset are used, as these two datasets are widely employed in existing methods. CIFAR-100 is composed of $32\times 32$ images taken from 100 classes, with 50,000/10,000 images for training/testing, and each class has the same number of training and testing images. Cityscapes is designed for understanding urban scenes and includes 2,975/500/1,525 images for training/validation/testing in 19 classes.
Table XII summarizes the distillation methods for classification and segmentation tasks. For each method, the knowledge source, teacher and student architectures, and accuracy (for classification) and mIoU (for segmentation) after distillation are reported. For a fair comparison, numbers are directly reported from the corresponding paper. However, in some cases where the original paper does not provide results on the mentioned datasets, and direct comparison is not possible, results obtained from implementing their models are reported (denoted by *).
In a similar vein, Table XIII presents a comparison of distillation methods for LLMs. It shows the compression rate and the improvement over the teacher model for each method. As can be seen, distillation in LLMs is more crucial than in other fields, as it allows for compressing large models with a high compression rate while still maintaining comparable performance. It should be noted that in cases where the exact performance of the teacher model is not reported, performance of the student model before and after distillation is reported (denoted by *).
VIII Conclusion and Discussion
@@
In the following, the current challenges of existing distillation methods are discussed, and insights into future research directions in knowledge distillation are provided.
VIII-A Challenges
Knowledge Extraction: Finding an appropriate source of knowledge for distillation is one of the main challenges in the distillation process [66, 791, 792]. Although using logits is the most straightforward approach to distillation, its effectiveness varies across different scenarios. For example, in tasks where labels are not available, logit-based distillation is not feasible. Furthermore, with the emergence of foundation models like CLIP, distilling rich features and embeddings has become crucial, which is not achievable through logit-based distillation alone, requiring feature-based and similarity-based distillation methods. In LLMs, this issue is even more pronounced, as some recent LLMs are not open-source [487, 488], restricting access to intermediate features or logits, and making black-box distillation particularly challenging. Existing methods often leverage a combination of different distillation sources to effectively transfer the teacher model’s knowledge to the student model.
Choosing Proper Distillation Scheme: Selecting an appropriate teacher-student architecture is another challenge, as it depends on the specific context. In scenarios where a pre-trained large model is available, offline distillation tends to be more effective, whereas, in other cases, online distillation may prove to be a better alternative. In self-supervised learning, self-distillation schemes have shown to be highly effective. Additionally, using multiple teachers to train a smaller network is a promising approach, often referred to as a mixture of experts. However, this comes with its own challenges, including the need for more powerful memory and GPUs, as well as the complexity of loading different teachers onto separate resources. In most existing methods, offline distillation is more commonly used. Specifically, when distilling from foundation models or LLMs, training the teacher alongside the student is not feasible due to the immense number of parameters in the teacher model.
Capacity Gap: Another important challenge is the capacity gap between the teacher and student models. It may be assumed that a larger teacher always leads to higher performance improvement in the student network, but when the teacher and student networks have a considerable size difference, it can cause capacity gap issues [88, 29, 55]. This is because the receptive fields of the teacher and student models differ significantly. This challenge is even more pronounced in foundation models, as they contain a tremendous number of parameters, making the distillation of their knowledge into a smaller model challenging. Furthermore, in black-box distillation of LLMs, where intermediate layers are not accessible, directly transferring knowledge from the teacher to the student is a serious challenge. Existing methods employ various algorithms to reduce this gap and make the teacher’s knowledge more comprehensible for the student.
Architectural Difference: Differences in the architecture of the teacher and student models present another significant challenge, which can impact the performance of the distillation method. In scenarios where the teacher and student have different architectures, which is very common, transferring proper knowledge becomes a new challenge. This issue becomes apparent when distilling knowledge from a teacher to a student with a different architecture, often leading to performance degradation [88, 29]. In foundation models, the student’s architecture can differ significantly from that of the teacher’s. Although logit-based distillation remains effective in such scenarios, feature-based and similarity-based distillation methods face additional challenges. More specifically, selecting appropriate layers with similar semantic information for feature distillation plays an important role in the performance of the method, as the teacher and student may even have different inductive biases. In black-box distillation for LLMs, CoT distillation can help address this problem, as it does not force the student to mimic the exact intermediate layers of the teacher. Instead, it provides the student with intermediate reasoning steps derived from the teacher. However, generating proper reasoning and addressing the capacity gap between the teacher and student remain significant challenges.
VIII-B Future Directions
Feature-based Distillation: While logit-based and similarity-based distillation are effective, recent trends and results show that feature-based distillation can lead to even greater improvements in student performance. However, it can also be combined with other methods. Recent feature-based distillation approaches have demonstrated that instead of defining complex relationships, a simple transformation on the features of either the teacher or student is sufficient to achieve SOTA results in enhancing the student’s performance [71, 76]. Furthermore, with the rapid growth of foundation models like CLIP, distilling embeddings from these large-scale models is gaining increasing attention.
Adaptive Distillation: Despite the existence of numerous distillation approaches, effectively distilling informative knowledge from the teacher to the student remains largely unexplored [791, 792]. Several adaptive distillation methods have been proposed to effectively transfer the dark knowledge of the teacher while ignoring unused knowledge. This line of research is particularly interesting, as it prevents the distillation of incorrect knowledge to the student and helps reduce the gap between the teacher and student networks. This process closely resembles human learning: a teacher transfers knowledge to students, but directly transferring all knowledge is not always effective or practical [88, 18]. Therefore, various forms of adaptive distillation can be employed, such as using a teacher assistant, decreasing the impact of distillation toward the end of training, or applying adaptive loss functions based on the importance of samples. All of these approaches align with real-world learning scenarios in human training processes.
Distillation from Foundation Models: With the emergence of a variety of foundation and multi-modal models, they have become a rich source of knowledge for distillation [793]. For example, CLIP, with its image and text encoder, can provide student models with rich embeddings, where text can be used to improve vision tasks such as weakly-supervised semantic segmentation, making it a hot topic in research [533]. Additionally, due to the general-purpose features of these large-scale models, they can be leveraged for open-vocabulary tasks [604], where the task is not limited to predefined classes. These interesting applications of distillation present great potential for future research.
Distillation in LLMs: While distillation is a highly effective approach in vision, text, and speech tasks, its application in LLMs is even more important. As stated earlier, LLMs have been shown to be overparameterized [3, 4] with billions and even trillions of parameters, far exceeding the size of models in other fields. These weight matrices are often sparse [5], making distillation crucial in reducing model size while maintaining performance. Additionally, most powerful models are not open-source [487, 488], increasing the importance of black-box distillation. CoT distillation is a particularly interesting technique for transferring reasoning capabilities from an LLM to an SLM. Unlike conventional distillation methods, CoT aims to enhance the generalization ability of the student model by teaching it how to reason. Results show that by distilling the knowledge of an LLM, the student model can achieve performance comparable to the teacher while significantly reducing model size. This is especially important because the teacher model may be too large to run on common GPUs, whereas the distilled student model can be efficiently deployed on common hardware, making advanced AI more accessible to to a wider audience.
Applications of Distillation: In addition to proposing novel distillation methods, distillation has been applied across various tasks, resulting in significant performance improvements. For example, distillation has been successfully applied in adversarial attacks [794, 166], mitigating backdoor attacks [795, 796], and anomaly detection [797]. Furthermore, the concept of distillation is utilized in dataset distillation [798], making it an interesting research direction. Another application of distillation is its integration with other compression methods, such as quantization or low-rank factorization.

### v2

@@ 7 Performance Comparison
ILD (Huang et al., 2022e) ICL $GPT2_{large}$ 6.00 CrossFit 61.20 / 66.20 92.0%
LLM-R (Wang et al., 2023h) ICL $LLaMA_{13B}$ 2.00 CommonsenseQA 48.80 / 49.60 98.0%
AICD (Liu, 2024) ICL GPT-3.5-Turbo SVAMP 51.70 / 20.70 250.0%
UniversalNER (Zhou et al., 2023c) IF GPT-3.5-Turbo UNIVERSAL NER 41.70 / 34.90 119.0% Link
LaMini-LM (Wu et al., 2023c) IF $Alpaca_{7B}$ 9.00 OpenBookQA 34.00 / 43.20 79.0% Link
Lion (Jiang et al., 2023b) IF $GPT_{175B}$ 25.00 BBH 32.00 / 48.90 65.0% Link
Fine-tune-CoT (Ho et al., 2022) CoT $GPT_{175B}$ 26.00 SVAMP 30.33 / 64.67 47.0% Link
Li et al. (Li et al., 2022e) CoT $GPT_{175B}$ 58.00 CommonsenseQA 82.47 / 73.00 113.0%
Magister et al. (Magister et al., 2022) CoT $GPT_{175B}$ 16.00 StrategyQA 63.77 / 65.40 97.0%
MCC-KD (Chen et al., 2023a) CoT GPT-3.5-Turbo SVAMP 68.66 / 75.14 91.0% Link
CSoTD (Li et al., 2023d) CoT $GPT_{175B}$ 135.00 CommonsenseQA 67.00 / 82.10 82.0%
SCOTT (Wang et al., 2023j) CoT $GPT\,neox_{20B}$ 7.00 CommonsenseQA 74.70 / 60.40 124.0% Link
Distilling step-by-step (Hsieh et al., 2023) CoT $Palm_{540B}$ 700.00 SVAMP 58.40 / 72.30 81.0% Link
PaD (Zhu et al., 2023b) CoT GPT-3.5-Turbo SVAMP 36.70 / 57.80 64.0%
TDG (Li et al., 2024g) CoT GPT-3.5-Turbo MATH 6.81 / 3.88 175.0% Link
RevThink (Chen et al., 2024c) CoT Gemini-1.5-Pro-001 30.00 CommonsenseQA 75.76 / 76.72 99.0%
8 Discussion
8.1 Principles of Method Classification
To structure this survey, the reviewed methods are categorized according to their primary contributions across five dimensions: distillation sources, schemes, algorithms, modalities, and applications. While many approaches span multiple categories, each method is assigned based on its central innovation, that is, the main idea or technical novelty the paper introduces, with relevant cross-references provided where appropriate. This principled organization aims to balance clarity with the multi-dimensional nature of knowledge distillation research.
8.2 Guidelines for Selecting Knowledge Distillation Strategies
To complement the presented classification, the survey also offers practical guidance for selecting appropriate KD strategies based on common task-specific constraints. While modality and application are typically determined by the nature of the target domain, the selection of distillation source, scheme, and algorithm remains a critical aspect of model design. Table 17 provides a decision-oriented summary to map task scenarios to suitable KD components. This perspective bridges the gap between methodological classification and practical usage, offering a starting point for design decisions in KD, while recognizing that the optimal choice ultimately depends on the specific goals, constraints, and context of each use case.
Table 17: Summary of key task constraints in knowledge distillation, with corresponding guidance for selecting suitable strategies, including KD sources, schemes, and suitable algorithms.
Task Requirement KD Source(s) KD Scheme(s) KD Algorithm(s)
Real-time Inference Logit-based Online Attention-based Adaptive
Data Scarcity / Few Shot Feature-based Self-distillation Graph-based Contrastive
Model Interpretability Feature-based Offline Attention-based Graph-based
Fine-grained Recognition Tasks Feature-based Similarity-based Offline Attention-based Contrastive
Limited Compute Budget Logit-based Offline Adaptive
Adversarial Robustness Logit-based Offline Adversarial
Multimodal Fusion Feature-based Offline Cross-modal
Continual/Incremental Learning Logit-based Similarity-based Online Adaptive
Multi-task Learning Logit-based Feature-based Online Adaptive Multi-teacher
Learning from Noisy Labels Similarity-based Self-distillation Contrastive
8.3 Quantitative Comparison
Quantitative comparison of existing distillation methods is very challenging. The performance of each method is influenced by numerous hyperparameters and is highly dependent on the weights of the pre-trained teacher model. As a result, providing a clear and fair comparison is nearly impossible, and reported performance metrics often differ significantly from those in the original papers. Therefore, we have intentionally avoided making direct claims about which method performs better than others.
However, in each section of the survey, we have highlighted the strengths and advantages of each category over others. In addition, Tables 12, 13, 14, 15, and 16 report the performance of various distillation methods across different domains, tasks, and datasets. Where available, we also include results for cases where the teacher and student models have either similar or different architectures. This information enables meaningful comparison, as it demonstrates the impact of varying dataset sizes and the robustness of each method to architectural differences.
8.4 Practical and Ethical Aspects
KD can also be studied from various other perspectives, including its use cases beyond model compression, such as enhancing privacy, improving security, and addressing intellectual property concerns. While this is beyond the scope of this paper, it is important to highlight some of the key challenges in these areas.
Recent studies have explored the use of KD to enhance privacy and prevent data leakage. For instance, Fernandez et al. (2023) employed KD to mitigate data memorization in text-to-image generative models, Flemings & Annavaram (2024) applied it to reduce the risk of LLMs generating sensitive information, and Shao et al. (2024a) utilized KD to train a large teacher model without exposing the data of smaller edge devices. On the other hand, Zhao & Zhang (2025) demonstrated that techniques like Data-Free Knowledge Distillation (DFKD) do not inherently guarantee privacy, and Cui et al. (2025) showed that such approaches may even introduce new vulnerabilities in LLMs. KD can also be used to guide the student model—intended as the final model—toward more fair and equitable behavior. (Chai et al., 2022)
Another important real-world consideration of KD is its potential to reduce computational demands, particularly relevant in the context of LLMs, which typically require substantial processing power. By enabling student models to achieve comparable performance with significantly lower resource consumption, KD can contribute to more energy-efficient and environmentally sustainable AI systems (Yuan et al., 2024c). However, the use of commercial models as teachers in the distillation process may raise legal concerns, as it could violate copyright protections. This limitation must be carefully considered in both research and deployment contexts.
8.5 Challenges
Knowledge Extraction: Finding an appropriate source of knowledge for distillation is one of the main challenges in the distillation process (Heo et al., 2019a; Hsu et al., 2022; Ojha et al., 2023). Although using logits is the most straightforward approach to distillation, its effectiveness varies across different scenarios. For example, in tasks where labels are not available, logit-based distillation is not feasible. Furthermore, with the emergence of foundation models like CLIP, distilling rich features and embeddings has become crucial, which is not achievable through logit-based distillation alone, requiring feature-based and similarity-based distillation methods. In LLMs, this issue is even more pronounced, as some recent LLMs are not open-source (Achiam et al., 2023; Team et al., 2023), restricting access to intermediate features or logits, and making black-box distillation particularly challenging. Existing methods often leverage a combination of different distillation sources to effectively transfer the teacher model’s knowledge to the student model.
Choosing Proper Distillation Scheme: Selecting an appropriate teacher-student architecture is another challenge, as it depends on the specific context. In scenarios where a pre-trained large model is available, offline distillation tends to be more effective, whereas, in other cases, online distillation may prove to be a better alternative. In self-supervised learning, self-distillation schemes have shown to be highly effective. Additionally, using multiple teachers to train a smaller network is a promising approach, often referred to as a mixture of experts. However, this comes with its own challenges, including the need for more powerful memory and GPUs, as well as the complexity of loading different teachers onto separate resources. In most existing methods, offline distillation is more commonly used. Specifically, when distilling from foundation models or LLMs, training the teacher alongside the student is not feasible due to the immense number of parameters in the teacher model. Table 17 presents recommended distillation strategies tailored to specific conditions.
Capacity Gap: Another important challenge is the capacity gap between the teacher and student models. It may be assumed that a larger teacher always leads to higher performance improvement in the student network, but when the teacher and student networks have a considerable size difference, it can cause capacity gap issues (Cho & Hariharan, 2019; Guo et al., 2020a; Huang et al., 2022a). This is because the receptive fields of the teacher and student models differ significantly. This challenge is even more pronounced in foundation models, as they contain a tremendous number of parameters, making the distillation of their knowledge into a smaller model challenging. Furthermore, in black-box distillation of LLMs, where intermediate layers are not accessible, directly transferring knowledge from the teacher to the student is a serious challenge. Existing methods employ various algorithms to reduce this gap and make the teacher’s knowledge more comprehensible for the student.
Architectural Difference: Differences in the architecture of the teacher and student models present another significant challenge, which can impact the performance of the distillation method. In scenarios where the teacher and student have different architectures, which is very common, transferring proper knowledge becomes a new challenge. This issue becomes apparent when distilling knowledge from a teacher to a student with a different architecture, often leading to performance degradation (Cho & Hariharan, 2019; Guo et al., 2020a). In foundation models, the student’s architecture can differ significantly from that of the teacher’s. Although logit-based distillation remains effective in such scenarios, feature-based and similarity-based distillation methods face additional challenges. More specifically, selecting appropriate layers with similar semantic information for feature distillation plays an important role in the performance of the method, as the teacher and student may even have different inductive biases. In black-box distillation for LLMs, CoT distillation can help address this problem, as it does not force the student to mimic the exact intermediate layers of the teacher. Instead, it provides the student with intermediate reasoning steps derived from the teacher. However, generating proper reasoning and addressing the capacity gap between the teacher and student remain significant challenges. Tables 12, 13, 14, and 15 present the effect of teacher-student architectural differences on the performance of knowledge distillation methods.
8.6 Future Directions
Feature-based Distillation: While logit-based and similarity-based distillation are effective, recent trends and results show that feature-based distillation can lead to even greater improvements in student performance. However, it can also be combined with other methods. Recent feature-based distillation approaches have demonstrated that instead of defining complex relationships, a simple transformation on the features of either the teacher or student is sufficient to achieve SOTA results in enhancing the student’s performance (Liu et al., 2023c; 2024g). Furthermore, with the rapid growth of foundation models like CLIP, distilling embeddings from these large-scale models is gaining increasing attention.
Adaptive Distillation: Despite the existence of numerous distillation approaches, effectively distilling informative knowledge from the teacher to the student remains largely unexplored (Hsu et al., 2022; Ojha et al., 2023). Several adaptive distillation methods have been proposed to effectively transfer the dark knowledge of the teacher while ignoring unused knowledge. This line of research is particularly interesting, as it prevents the distillation of incorrect knowledge to the student and helps reduce the gap between the teacher and student networks. This process closely resembles human learning: a teacher transfers knowledge to students, but directly transferring all knowledge is not always effective or practical (Cho & Hariharan, 2019; Mirzadeh et al., 2020). Therefore, various forms of adaptive distillation can be employed, such as using a teacher assistant, decreasing the impact of distillation toward the end of training, or applying adaptive loss functions based on the importance of samples. All of these approaches align with real-world learning scenarios in human training processes.
Distillation from Foundation Models: With the emergence of a variety of foundation and multi-modal models, they have become a rich source of knowledge for distillation (Zhang et al., 2024d). For example, CLIP, with its image and text encoder, can provide student models with rich embeddings, where text can be used to improve vision tasks such as weakly-supervised semantic segmentation, making it a hot topic in research (Xie et al., 2022). Additionally, due to the general-purpose features of these large-scale models, they can be leveraged for open-vocabulary tasks (Liu et al., 2024f), where the task is not limited to predefined classes. These interesting applications of distillation present great potential for future research.
Distillation in LLMs: While distillation is a highly effective approach in vision, text, and speech tasks, its application in LLMs is even more important. As stated earlier, LLMs have been shown to be overparameterized (Kaplan et al., 2020; Fischer et al., 2024) with billions and even trillions of parameters, far exceeding the size of models in other fields. These weight matrices are often sparse (Hu et al., 2021), making distillation crucial in reducing model size while maintaining performance. Additionally, most powerful models are not open-source (Achiam et al., 2023; Team et al., 2023), increasing the importance of black-box distillation. CoT distillation is a particularly interesting technique for transferring reasoning capabilities from an LLM to an SLM. Unlike conventional distillation methods, CoT aims to enhance the generalization ability of the student model by teaching it how to reason. Results show that by distilling the knowledge of an LLM, the student model can achieve performance comparable to the teacher while significantly reducing model size. This is especially important because the teacher model may be too large to run on common GPUs, whereas the distilled student model can be efficiently deployed on common hardware, making advanced AI more accessible to to a wider audience.
Besides its role in training SLMs, a promising new direction is the use of KD during the pre-training stage of LLMs in industry (Yang et al., 2025; Team et al., 2025). Training LLMs of different sizes is highly cost-intensive. KD plays an important role in creating smaller versions of a pre-trained LLM, where techniques such as quantization (Liu et al., 2023b) or pruning (Sreenivas et al., 2024; Muralidharan et al., 2024) are applied to the large model to extract a compact student model. This process can be used in a post-training setting, requiring only a small fraction of the dataset originally used to train the teacher model.
Applications of Distillation: In addition to proposing novel distillation methods, distillation has been applied across various tasks, resulting in significant performance improvements. For example, distillation has been successfully applied in adversarial attacks (Papernot et al., 2016b; Goldblum et al., 2020), mitigating backdoor attacks (Li et al., 2021a; Han et al., 2024b), and anomaly detection (Salehi et al., 2021). Furthermore, the concept of distillation is utilized in dataset distillation (Wang et al., 2018a), making it an interesting research direction. Another application of distillation is its integration with other compression methods, such as quantization or low-rank factorization.
9 Conclusion